19. A parallel plate capacitor has a capacitance of 600F when the plate separation is 6mm. If the plate separation is changed to 5mm, the new value of the capacitance is

Group of answer choices

450F.

500F.

675F.

720F.

900F.
Answer #1

Capacitance= AEo/d

A-> area

d= separation between plates

C1/C2= d2/d1

600/C2= 5/6

C2= 720 F

Option 4 is correct.